OpenStudy (anonymous):

#10 Solve the problem using the unpaid balance method. The account balance on April 1st is $60.15. On April 15th a payment of $51.00 is made. On April 25th a purchase of $91.27 is made. What is the finance charge if the annual rate is 18%? What is the new account balance? Finance charge = $______ New account balance = $_____

OpenStudy (texaschic101):

60.15 - 51.00 = 9.15 (balance - payment = new balance) 9.15 + 91.27 = 100.42 (balance + purchase = new balance, not including finance charge Now we need to figure the finance charge... 18% annually = 18%/12 = 1.5% per month 100.42 x 1.5% = 100.42(.015) = 1.51 < -- finance charge 100.42 + 1.51 = 101.93 < -- new balance

OpenStudy (texaschic101):

@whpalmer4 ...can you check this for me....please :)

OpenStudy (texaschic101):

or do I need to do finance charge on 60.15 and add that to the problem ?

OpenStudy (texaschic101):

@mathstudent55 ....can you take a peek at this :)

OpenStudy (texaschic101):

@satellite73 ....help please

OpenStudy (texaschic101):

do I need to do the finance charge on the 60.15 and not the 100.42 ?

OpenStudy (texaschic101):

or both ?

OpenStudy (texaschic101):

@ConDawg ...can you take a peek at this

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

hmm.. So there's a balance of $60.15, and then the makes a payment with that money so that's taking away from the balance, taking away 51$ But the it says she made a purchase of $91.27 that sounds like to me that it is also taking from the balance, but that doesn't make much sense...

OpenStudy (texaschic101):

a purchase of 91.27 was made and that adds to the balance.....I just dont know about the finance charge.....the 60.15 is the unpaid balance from the previous month....does the finance charge go on that and the 100.42 or just the 60.15 ?

OpenStudy (texaschic101):

Thinking about it ....maybe it should go on both

OpenStudy (texaschic101):

@phi ...can you help us

OpenStudy (anonymous):

We have a time frame of 25 day's right? We also have an annual rate of 18% per year. That means it take 18% of what you have at the end of the year. We are in the time frame of less than a month. So 18/12=1.5% per month. If we want to get picky, there are 30 days in April and we are talking about 25 of them. So 25/30=5/6, 5/6* 1.5 = 5/4% so we want the 5/4% at the end of 25 days, which is 100.42. Anyways, that's the way I take it. This is why I hate word problems like this! It becomes more about the English and what's being said than the actual math! There's a difference between a lawyer and a mathematician.

OpenStudy (texaschic101):

lol....you are so right.......so you don't think there is a finance charge on the 60.15 ?

OpenStudy (texaschic101):

It is the unpaid balance from the previous month

OpenStudy (texaschic101):

@ganeshie8 .....please help

OpenStudy (anonymous):

I do not think the finance charge in on the 60.15, but IS on the 100.42

OpenStudy (texaschic101):

that is what I thought at first, but now I am questioning myself

OpenStudy (anonymous):

Yeah I know what you mean.

OpenStudy (texaschic101):

@phi .....do you know ?

OpenStudy (phi):

You do not pay a finance charge on the $91.27 if you buy something on a credit card, it shows up on the next bill. If you pay it off in total, there is no finance charge. I think you either pay interest on $60.15 or the difference between 60.15 and 51.00 Not sure, but I think it is the first.

OpenStudy (texaschic101):

oh.....so pay finance charge on just the 60.15 ?

OpenStudy (phi):

It depends if 60.15 was left over from the previous bill (in which case you pay interest), or it was charged last month, and you are making a partial payment, leaving 9.15 unpaid (so interest will be added)

OpenStudy (texaschic101):

The question is an unpaid balance so I assumed that the 60.15 was the unpaid balance.....I guess that is what I get for assuming

OpenStudy (texaschic101):

I am sorry but I got to go now......thanks guys, you are great

OpenStudy (whpalmer4):

@phi you don't pay any finance charge only if you had previously paid off your entire balance. if even 1 penny of your previous balance carried over, then your new purchases accrue interest throughout the billing period, starting on the date of purchase. typically the average daily balance is used, and the interest rate is 1/365th of the nominal rate, compounded daily. @texaschic101 Now, this problem says it uses the unpaid balance method, which I haven't personally encountered, but google is our friend, and to compute interest for the UBM, you take the starting balance ($60.15 here), compute interest on that for the month, then add the net of any purchases and payments made during the month, and that's your new balance. Typical billing cycles are around 25 days, so it is reasonable IMO to assume the balance of $60.15 on April 1 is in fact the unpaid balance from a prior billing cycle given that the end of the problem is no earlier than April 25. Also, if it isn't a prior balance, then why would we be asked to compute the interest, as there wouldn't be any with the unpaid balance method if the prior balance was $0.

OpenStudy (texaschic101):

so would it be : 60.51 x 1.5% = 60.51(.015) = 0.91 <--finance charge 60.51 + 0.91 = 61.42 61.42 - 51.00 = 10.42 (subtract 51.00 payment 10.42 + 91.27 = 101.69 (add 91.27 purchase)..and this is your new balance

OpenStudy (whpalmer4):

Now I've found some other websites that suggest the procedure is to take the previous closing balance ($60.15), subtract any payments or credits, and that is the unpaid balance. Compute the interest on that amount, then add the interest and any purchases to arrive at the new balance. To put it plainly, you pay interest on the portion of the previous balance which hasn't been paid off by the end of the billing cycle.